OpenStudy (anonymous):

find the slope of the line containing the pair of points (0,-7) and (1,-2)

OpenStudy (chaise):

Use the point slope formula and also the gradient formula y=mx+c \[m=\frac{y-y}{x-x}\]

OpenStudy (anonymous):

can you show me how i am not sure

OpenStudy (anonymous):

(0,-7) and (1,-2) \[ \frac { -2 -(-7)}{1-0}=? \]

OpenStudy (anonymous):

slope would be 5 then right?

OpenStudy (chaise):

yes, now you have a point (x,y) = (1, -2) and a slope (m) plug it into this equation and find c. y=mx+c

OpenStudy (anonymous):

why would i do that i just needed to find the slope and the slope is 5 i just wanted someone to double check for me to make sure i am correct